# Linear-Size $\mathrm{QAC}^{0}$ Channels: Learning, Testing and Hardness

### v1

@@ Acknowledgement
This work was supported by National Natural Science Foundation of China (Grant No. 62332009, and 12347104), Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302901), NSFC/RGC Joint Research Scheme (Grant No. 12461160276), Natural Science Foundation of Jiangsu Province (Grant No. BK20243060) and the New Cornerstone Science Foundation.
@@ 1.3 Summary and Future Work
    While learning quantum states prepared by shallow circuits has been studied extensively in the PAC setting, little is known about agnostic learning in this context. Recently, the authors in [BBK+25] proposed an agnostic learning algorithm for product states. A natural question is whether similar techniques can be extended to efficiently learn states generated by $\textsf{state-}\mathbf{QNC}^{0}$ or $\textsf{state-}\mathbf{QLC}^{0}$?
@@ 2 Preliminaries
Proposition 2.3 ([Bha13, Theorem X.1.1]).
@@
Proposition 2.5 ([HKOT23, Proposition 1.6]).
@@
For further details regarding distance relations between quantum channels, readers may refer to Yuan and Fung’s work [YF17].
@@ 2.1 Analysis of Boolean Functions
In this subsection, we briefly introduce the theory of analysis of Boolean functions. Readers may refer to O’Donnell’s excellent book [O’D14] for a thorough treatment.
@@ 2.3 Agnostic learning
Kearns, Schapire and Sellie [KSS92] proposed the agnostic PAC learning model, which is a more general learning model, capturing the case where the learning object may not be in the hypothesis class.
@@
The agnostic model is believed to be closer to the realistic scenario than the standard PAC model, especially in quantum computing, where noise is unavoidable. However, designing efficient agnostic learning algorithms is generally challenging. Even very few concept classes are known to be agnostically learnable in subexponential time. Bun, Kothari, and Thaler [BKT19] gave a subexponential time agnostic learning algorithm for the class of functions with approximate degree $n^{c}$ for $c<1$, built on [KKMS08].
@@
    Quantum Process Statistical Query (QPSQ) Model [WD25]. For an $n\to m$ quantum channel $\Phi$, in one query we can only obtain classical information about $\Phi$. That is, we can access $\Phi$ by first preparing a quantum state $\varphi$ on $n$ qubits, applying the quantum channel to get the quantum state $\Phi(\varphi)$, and immediately applying a quantum measurement.
In this work, we consider the agnostic learning of quantum channels, which is also referred to as Agnostic Process Tomography introduced by Wadhwa, Lewis, Kashefi and Doosti [WLKD24].
@@ 2.5 Quantum circuit
In this work, we are concerned with $\mathbf{QAC}^{0}$ circuits, which are polynomial-size constant-depth quantum circuits consisting of single-qubit unitaries and multi-qubit CZ-gates [^2] Some definitions use generalized Toffoli gates. They are equivalent (by the reduction in [FFG+06]) in our case.. An $n$-qubit CZ-gate is defined as
@@
We also introduce the class of $\mathbf{QLC}^{0}$ circuits, which consists of $\mathbf{QAC}^{0}$ circuits with linear-size ancilla. $\mathbf{QLC}^{0}$ is a quantum counterpart of the classical circuit family $\mathbf{LC}^{0}$, introduced by Chaudhuri and Radhakrishnan [CR96], which is one of the most interesting subclasses of $\mathbf{AC}^{0}$ and has received significant attention from various perspectives [KPLT06, CIP09, SS12]. A quantum channel induced by a $\mathbf{QLC}^{0}$ circuit is defined analogously.
@@ 3 Approximation of $\mathbf{QAC}^{0}$ circuits
In this section, we use unitary dilation combined with the low-degree approximation for CZ-gates, to prove a new low-degree approximation for $\mathbf{QAC}^{0}$ circuits, improving the results of Anshu, Dong, Ou, and Yao [ADOY25].
@@
This result is incomparable to [ADOY25]. Although we achieve a better approximation degree upper bound, we have the restriction in Eq. 7 for the error parameter $r$ in this work.
Note that when the degree of $A$ is upper bounded by $\ell=n^{o(1)}$, the above theorem holds for any $r=\operatorname{polylog}(n)$. So we can recover all the results in [ADOY25] with better approximate degree upper bound parameters. For example, Theorem 3.1 implies that any $\mathbf{QAC}^{0}$ circuit that computes $\operatorname{Parity}_{n}$ requires at least $a=\tilde{\Omega}\left(n^{1+2^{-d}}\right)$ ancilla, slightly improving the $\tilde{\Omega}\left(n^{1+3^{-d}}\right)$ lower bound in [ADOY25].
@@ 3.1 Unitary Dilation
We start with the unitary dilation technique used in this work, which is inspired by Montanaro, Shao, and Verdon [MSV24] [^3] This work is withdrawn by the authors due to some flaws in their proof.. The unitary dilation technique allows us to treat non-unitary operators as unitaries by embedding them into larger unitary operators. This is especially useful when applying the polynomial methods [Smo87, Tar92] to quantum circuits. In particular, the low-degree approximations of large quantum gates are not necessarily unitary operators. By embedding them into larger unitary operators, we obtain a unitary approximation with a low-degree submatrix.
@@ 3.2 Proofs of Low-Degree Approximation
Lemma 3.9 (Approximation for CZ-gate, [ADOY25, Corollary 3.3]).
@@
We note that, in contrast to [ADOY25], our approximation requires a range constraint on the parameter $r$. Fortunately, if we require an inverse-polynomial error, we only need $r=O(\log n)$ and this remains within the applicable scope of the lemma.
@@ 4 Agnostic Learning for $\mathbf{QLC}^{0}$ channels
In this section, we present an agnostic learning algorithm for $\mathbf{QLC}^{0}$ channels with respect to the Frobenius norm as in Algorithm 1. This algorithm employs the classical shadow tomography introduced by Huang, Kueng and Preskill [HKP20].
Lemma 4.1 (Classical Shadows for Low-degree Pauli Operators, [HKP20, Lemma 17]).
@@ 4.1 Analysis of Agnostic Learning Algorithm
To prove 4.7, this work requires some results in [ADOY25]. We therefore reformulate the results accordingly.
Lemma 4.8 (Approximate degree for state, [AM23], [ADOY25, Corollary 3.2]).
@@
Lemma 4.10 ([ADOY25, Lemma 2.12]).
@@
With [ADOY25, Fact 7.1],
@@
The proof of this proposition is highly similar to that of 4.7, and the argument is completed by simply replacing Theorem 3.1 with [ADOY25, Lemma 3.5].
@@ 6.1 Hardness Reduction
Unlike [VH25] and [HLB+24], we perform no operations during step B of our protocol.
@@
Let us focus on the step C. We review some facts originating from quantum information theory (see the book of Watrous [Wat18] for more information): if an $n\to 1$ channel $\Phi$ has the form $\rho\to\mathrm{Tr}_{-i}\>\!\!\left(V\rho V^{\dagger}\right)$, then its dual channel $\Phi^{*}$ will take the form $\rho_{i}\to V^{\dagger}(\rho_{i}\otimes\mathds{1}_{-i})V$. Moreover, the Choi representation of the dual channel and that of the original channel satisfy the following relation:
@@ 6.2 Hardness Results
Let us briefly recall the background of the reduction, where we take the same setting as in Vasconcelos and Huang’s work [VH25]. Suppose $U$ is a unitary generated by a circuit with $(n+a)$ qubits. We only consider circuits where the ancilla qubits are initialized to the $\left|0_{a}\right\rangle$ state and the computation is clean. Since the computation is clean, $U$ implicitly characterizes a unitary matrix $V$ acting on $n$ input qubits.
@@
Now we are ready to prove Theorem 6.1. Let us first examine the learning hardness of $\mathbf{QAC}^{0}$ unitary. We use a result of Vasconcelos and Huang [VH25], tailored to our needs. For completeness, we provide a proof.
Lemma 6.7 (Tailored from [VH25, Propsition 7]).
@@
$U_{x}$ is in the depth-$d$ $\mathbf{QAC}^{0}$ circuit with ancilla even for $a=0$ and $d=1$. However, distinguishing $\mathds{1}$ from one of $U_{x}$ has a Grover Search lower bound from Bennett, Bernstein, Brassard and Vazirani’s work [BBBV97] with queries at least $\exp\left(\Omega(n)\right)$. This also means distinguishing $\mathds{1}$ from one of $U_{x}\otimes U_{x}^{\dagger}$ has an exponential bound. ∎
@@ Appendix A Proofs
The proof is similar to the one in [ADOY25].

### v2

@@ Acknowledgement
This work was supported by National Natural Science Foundation of China (Grant No. 62332009, and 12347104), Quantum Science and Technology-National Science and Technology Major Project (Grant No. 2021ZD0302901), NSFC/RGC Joint Research Scheme (Grant No. 12461160276), Natural Science Foundation of Jiangsu Province (Grant No. BK20243060) and the New Cornerstone Science Foundation.
@@ 1.3 Summary and Future Work
    While learning quantum states prepared by shallow circuits has been studied extensively in the PAC setting, little is known about agnostic learning in this context. Recently, the authors in [10.1145/3717823.3718207] proposed an agnostic learning algorithm for product states. A natural question is whether similar techniques can be extended to efficiently learn states generated by $\textsf{state-}\mathbf{QNC}^{0}$ or $\textsf{state-}\mathbf{QLC}^{0}$?
@@ 2 Preliminaries
Proposition 2.3 ([bhatia2013matrix, Theorem X.1.1]).
@@
Proposition 2.5 ([Haah_2023, Proposition 1.6]).
@@
For further details regarding distance relations between quantum channels, readers may refer to Yuan and Fung’s work [Yuan_2017].
@@ 2.1 Analysis of Boolean Functions
In this subsection, we briefly introduce the theory of analysis of Boolean functions. Readers may refer to O’Donnell’s excellent book [ODonnell2014] for a thorough treatment.
@@ 2.3 Agnostic learning
Kearns, Schapire and Sellie [10.1145/130385.130424] proposed the agnostic PAC learning model, which is a more general learning model, capturing the case where the learning object may not be in the hypothesis class.
@@
The agnostic model is believed to be closer to the realistic scenario than the standard PAC model, especially in quantum computing, where noise is unavoidable. However, designing efficient agnostic learning algorithms is generally challenging. Even very few concept classes are known to be agnostically learnable in subexponential time. Bun, Kothari, and Thaler [doi:10.1137/1.9781611975482.42] gave a subexponential time agnostic learning algorithm for the class of functions with approximate degree $n^{c}$ for $c<1$, built on [doi:10.1137/060649057].
@@
    Quantum Process Statistical Query (QPSQ) Model [Wadhwa_2025]. For an $n\to m$ quantum channel $\Phi$, in one query we can only obtain classical information about $\Phi$. That is, we can access $\Phi$ by first preparing a quantum state $\varphi$ on $n$ qubits, applying the quantum channel to get the quantum state $\Phi(\varphi)$, and immediately applying a quantum measurement.
In this work, we consider the agnostic learning of quantum channels, which is also referred to as Agnostic Process Tomography introduced by Wadhwa, Lewis, Kashefi and Doosti [wadhwa2024agnostic].
@@ 2.5 Quantum circuit
In this work, we are concerned with $\mathbf{QAC}^{0}$ circuits, which are polynomial-size constant-depth quantum circuits consisting of single-qubit unitaries and multi-qubit CZ-gates [^2] Some definitions use generalized Toffoli gates. They are equivalent (by the reduction in [10.5555/2011679.2011682]) in our case.. An $n$-qubit CZ-gate is defined as
@@
We also introduce the class of $\mathbf{QLC}^{0}$ circuits, which consists of $\mathbf{QAC}^{0}$ circuits with linear-size ancilla. $\mathbf{QLC}^{0}$ is a quantum counterpart of the classical circuit family $\mathbf{LC}^{0}$, introduced by Chaudhuri and Radhakrishnan [10.1145/237814.237824], which is one of the most interesting subclasses of $\mathbf{AC}^{0}$ and has received significant attention from various perspectives [1663737, 10.1007/978-3-642-11269-0_6, 10.1007/978-3-642-31594-7_65]. A quantum channel induced by a $\mathbf{QLC}^{0}$ circuit is defined analogously.
@@ 3 Approximation of $\mathbf{QAC}^{0}$ circuits
In this section, we use unitary dilation combined with the low-degree approximation for CZ-gates, to prove a new low-degree approximation for $\mathbf{QAC}^{0}$ circuits, improving the results of Anshu, Dong, Ou, and Yao [ADOY24].
@@
This result is incomparable to [ADOY24]. Although we achieve a better approximation degree upper bound, we have the restriction in Eq. 7 for the error parameter $r$ in this work.
Note that when the degree of $A$ is upper bounded by $\ell=n^{o(1)}$, the above theorem holds for any $r=\operatorname{polylog}(n)$. So we can recover all the results in [ADOY24] with better approximate degree upper bound parameters. For example, Theorem 3.1 implies that any $\mathbf{QAC}^{0}$ circuit that computes $\operatorname{Parity}_{n}$ requires at least $a=\tilde{\Omega}\left(n^{1+2^{-d}}\right)$ ancilla, slightly improving the $\tilde{\Omega}\left(n^{1+3^{-d}}\right)$ lower bound in [ADOY24].
@@ 3.1 Unitary Dilation
We start with the unitary dilation technique used in this work, which is inspired by Montanaro, Shao, and Verdon [montanaro2024lowdegreeapproximationqac0circuits] [^3] This work is withdrawn by the authors due to some flaws in their proof.. The unitary dilation technique allows us to treat non-unitary operators as unitaries by embedding them into larger unitary operators. This is especially useful when applying the polynomial methods [10.1145/28395.28404, tarui1992low] to quantum circuits. In particular, the low-degree approximations of large quantum gates are not necessarily unitary operators. By embedding them into larger unitary operators, we obtain a unitary approximation with a low-degree submatrix.
@@ 3.2 Proofs of Low-Degree Approximation
Lemma 3.9 (Approximation for CZ-gate, [ADOY24, Corollary 3.3]).
@@
We note that, in contrast to [ADOY24], our approximation requires a range constraint on the parameter $r$. Fortunately, if we require an inverse-polynomial error, we only need $r=O(\log n)$ and this remains within the applicable scope of the lemma.
@@ 4 Agnostic Learning for $\mathbf{QLC}^{0}$ channels
In this section, we present an agnostic learning algorithm for $\mathbf{QLC}^{0}$ channels with respect to the Frobenius norm as in Algorithm 1. This algorithm employs the classical shadow tomography introduced by Huang, Kueng and Preskill [Huang_2020].
Lemma 4.1 (Classical Shadows for Low-degree Pauli Operators, [Huang_2020, Lemma 17]).
@@ 4.1 Analysis of Agnostic Learning Algorithm
To prove 4.7, this work requires some results in [ADOY24]. We therefore reformulate the results accordingly.
Lemma 4.8 (Approximate degree for state, [Anshu2023concentrationbounds], [ADOY24, Corollary 3.2]).
@@
Lemma 4.10 ([ADOY24, Lemma 2.12]).
@@
With [ADOY24, Fact 7.1],
@@
The proof of this proposition is highly similar to that of 4.7, and the argument is completed by simply replacing Theorem 3.1 with [ADOY24, Lemma 3.5].
@@ 6.1 Hardness Reduction
Unlike [vasconcelos2024learning] and [HLB+24], we perform no operations during step B of our protocol.
@@
Let us focus on the step C. We review some facts originating from quantum information theory (see the book of Watrous [watrous2018theory] for more information): if an $n\to 1$ channel $\Phi$ has the form $\rho\to\mathrm{Tr}_{-i}\>\!\!\left(V\rho V^{\dagger}\right)$, then its dual channel $\Phi^{*}$ will take the form $\rho_{i}\to V^{\dagger}(\rho_{i}\otimes\mathds{1}_{-i})V$. Moreover, the Choi representation of the dual channel and that of the original channel satisfy the following relation:
@@ 6.2 Hardness Results
Let us briefly recall the background of the reduction, where we take the same setting as in Vasconcelos and Huang’s work [vasconcelos2024learning]. Suppose $U$ is a unitary generated by a circuit with $(n+a)$ qubits. We only consider circuits where the ancilla qubits are initialized to the $\left|0_{a}\right\rangle$ state and the computation is clean. Since the computation is clean, $U$ implicitly characterizes a unitary matrix $V$ acting on $n$ input qubits.
@@
Now we are ready to prove Theorem 6.1. Let us first examine the learning hardness of $\mathbf{QAC}^{0}$ unitary. We use a result of Vasconcelos and Huang [vasconcelos2024learning], tailored to our needs. For completeness, we provide a proof.
Lemma 6.7 (Tailored from [vasconcelos2024learning, Propsition 7]).
@@
$U_{x}$ is in the depth-$d$ $\mathbf{QAC}^{0}$ circuit with ancilla even for $a=0$ and $d=1$. However, distinguishing $\mathds{1}$ from one of $U_{x}$ has a Grover Search lower bound from Bennett, Bernstein, Brassard and Vazirani’s work [doi:10.1137/S0097539796300933] with queries at least $\exp\left(\Omega(n)\right)$. This also means distinguishing $\mathds{1}$ from one of $U_{x}\otimes U_{x}^{\dagger}$ has an exponential bound. ∎
@@ Appendix A Proofs
The proof is similar to the one in [ADOY24].